# Referential communication in heterogeneous communities of pre-trained visual deep networks

### v5

@@ 4.4.2 Disentangling meaning with a Sparse Autoencoder
To investigate which high-level features are used in communication, we use dictionary learning, which has proven useful to identify interpretable features in multimodal models (Parekh et al., 2024), and more specifically Sparse AutoEncoders (SAEs) (Bricken et al., 2023; Cunningham et al., 2023; Rajamanoharan et al., 2024; Templeton et al., 2024). In our case, the SAE consists of single-layer encoders and decoders, with an added ReLu nonlinearity for the encoder. The encoder projects 64-dimensional messages to an expanded hidden space (we use a hidden space of $16\times{}64=1024$ dimensions). Training is done with a mean-square error reconstruction loss, and an L1 regularization loss to add sparsity (weight of the L1 term is set to 0.01). The sparsity component of the loss ensures that meaningful information is stored in individual dimensions. We train the SAE using messages corresponding to images from the communication training set (ImageNet1k, see section 3.3). Informally varying the SAE hyperparameters did not impact our general conclusions. We present SAE results for messages from a VGG 11 sender + ViT-B/16 receiver pair, which are representative of other cases.
@@
First and most importantly, we observe that the dimensions uncovered by the SAE correspond to understandable semantic features. In particular, we qualitatively assess this by looking, for each SAE dimension, at the images corresponding to the messages closest to that dimension. [^9] We were able to do this for 1024 dimensions because many of them share the same neighbors. We observe that images clustered around a dimension share a common human-interpretable meaning. In Fig. 3, which contains the images closest to a representative SAE feature, we see different breeds of black dogs on white backgrounds. It is worth noting that they do not all share the same ImageNet1k class. The messages clustering around the same SAE vectors are abstract enough to overcome perceptual variation, with for example one of the images in Fig. 3 being a drawing (7th from the left), while the others are photographs. We present additional representative examples in Appendix M (Fig. 12). Additional evidence for the semantic nature of the messages is provided in a qualitative error analysis in Appendix N.
@@
SAE features also inform us about referent generalization. We use the same sender and SAE trained on ImageNet1k to extract the k-nearest messages from the OOD datasets to each one-hot SAE vector. In general, the closest OOD messages to each SAE vector also form human-interpretable clusters, noisily sharing properties with the corresponding ImageNet1k training images. In the case of Places205, bottom right in Fig. 3, white backgrounds (snow, table-cloth, floor) and dark objects (such as people wearing black clothing or killer whales) are present in most images, although none depicts dogs. For the more strongly OOD CelebA dataset, bottom left, pictures of celebrities tend to show either dark backgrounds, celebrities with dark skins, or in one case, makeup depicting an animal. This seems to capture only part of the properties associated to the SAE feature in domain. SAE features might be capturing combinations of primitive properties, such as foreground and background colours, that might only be partially shared by images that activate them in the message. Taking advantage of the continuous communication channel, messages can adjust distance to a given feature depending on the image. In OOD cases, where primitive properties might be less correlated than in the source domain, messages are further away from the one-hot feature than ImageNet1k images are, and only a subset of the primitive properties are captured. [^10] On average, there are 56 ImageNet1k messages closer to a feature than the closest CelebA message. For Places205, which is less OOD, there are on average 7 ImageNet1K messages closer to a feature.
@@
From another perspective, instead of looking at which OOD messages are nearer pre-selected SAE vectors, we can ask which SAE vectors are closer to any message from an OOD dataset. Messages describing OOD images cluster around SAE vectors with properties relevant to the dataset. The CelebA messages tend to be closer to the SAE dimensions representing people, while Places205 messages are closer to SAE vectors corresponding to urban settings, or landscapes. [^11] To avoid pareidolia, we verified this claim in a small experiment. We prepared 30 sets of 10 ImageNet1k images, corresponding either to a feature close to Places205 images, or to one feature corresponding to CelebA images, or to neither. One of the authors then labeled them, without knowing their source, for whether they contained foregrounded faces (close to CelebA), depicted landscapes or urban settings (close to Places205), or neither. Performance was well above random chance. We provide an example of each in Fig. 13 of Appendix M.
@@ 4.4.3 Perturbation analysis
In a next set of analyses, we gain insight on the information encoded in the message by looking at how image perturbations affect communication (an analysis of direct perturbation of the communication channel itself is reported in Appendix O). The effects of the perturbations on an example image are shown in Fig. 4. We perturb images from the ImageNet1k dataset that were seen by the agents during referential game training, where the receivers have perfect accuracy. In each experiment, we apply the same transformation to all images, both those seen by the sender and those seen by the receiver. Note that we test senders and receivers that have been trained with non-corrupted images, and only face the perturbations at test time.
Results for both the one-to-one and the population setups are reported in Fig. 5. All ablations affect accuracy. However, the color ablations (transforming the image to grayscale and jittering colors) and resizing the image are having a much smaller impact on performance than adding Gaussian blur. This is in line with our intuition about how much these changes should affect the ability to decode the contents of an image. As the example in Fig. 4 shows, color and size ablations don’t normally affect the human ability to discriminate the object depicted in the image, whereas blurring the image can easily make its contents difficult to decode for us as well.
We finally observe that all perturbation types affect one-to-one and population-trained agents in a similar manner, with population-trained agents being marginally more affected.
Figure 4: Perturbation examples
Figure 5: Accuracy after various image perturbations for the one-to-one and population setups. Gaussian blur was uniformly sampled within the [0.1,10] range. Mean accuracies and corresponding standard deviations are given across all 36 possible agent pairs. We add a small horizontal offset to the accuracy values for different perturbations to make them all visible. The horizontal lines connecting one-to-one and population accuracies for the same perturbation are meant to ease comparison between the two settings. Random baseline is 1.6% (randomly selecting one of the 64 images).
4.4.4 Specificities of population protocol
@@ 5 Conclusion
The networks can use the emergent protocol, to some extent, to communicate about referents that are only distinguishable at a more granular level than used during communication training; and they can, to some extent, communicate about referents coming from categories or even datasets that were not seen during pre-training nor in the communication protocol induction phase.
@@ Appendix A The one-to-one referential game in pseudocode
Appendix B Using visual representations without a communication layer
Our setup assumes that we need to connect networks that have already been pre-trained on visual tasks. While our main objective is to evaluate communication success, an interesting question is to what extent the communication layer is adding something non-trivial to what is already encoded in the pre-trained visual representations of the networks. To partially assess this, we play the game with “agents” that generate messages directly from last-layer visual representations. For each network, we clustered all images using k-means, setting $k$ to the ground-truth number of classes in the dataset. We then treated the k-means clusters as the “message” produced by the sender. The receiver would then randomly pick an image belonging to the correct cluster from the candidate set.
As table 5 shows, in the homogeneous case where sender and receiver have the same architecture, [^12] For the heterogeneous case, where sender and receiver are different networks, we tested various ways to align the clusters on the training set, but performance was consistently scarcely above chance. k-means pairs achieve on average ImageNet1k and OOD performance well above chance level, but clearly lagging behind the networks that underwent communication training (performance $>98\%$ and $>90\%$ on these tasks, respectively). This confirms that the latter has enriched the networks with discrimination abilities that go beyond what’s directly inherited from (mostly) supervised pre-training.
Table 5: Percentage accuracy of agents evaluated on the ImageNet1k and OOD sets. Agents are unsupervised clusters of visual representations built using the k-means algorithm.
ImageNet1k OOD
Homogeneous 49.8 $\pm$ 17.9 21.2 $\pm$ 7.3
@@ Appendix C Discrete experiments
In order to make sure that our negative results concerning discrete communication are not due to an unfortunate choice of hyperparameters, we repeated most experiments across a range of settings. We report here results for the more interesting population setup, but similar patterns are observed in one-to-one training. Table 8 shows that there is a small increase in accuracy when enlarging vocabulary size (compared to the results obtained with 256-dimensional messages reported in tables 6 and 7 above), but performance is still well below that of the continuous case even when using a vector as large as 8192 dimensions (compared to 16 or 64 dimensions in the continuous case). Note that using vocabulary sizes below 256 (not shown in the table) led to failure to converge to non-random accuracy.
@@
If using a discrete channel is considered a priority (for example, because it might lead to a more interpretable protocol), other discrete optimization techniques could also be explored, such as more advanced policy gradient methods (Francois-Lavet et al., 2018), or quantization approaches such as the one recently proposed by Carmeli et al. (2022) (although we tried something similar, with less-than-encouraging results, in Appendix O).
@@ Appendix G Augmentations and hard distractors
Results for agents trained with data augmentations are presented in Table 15. In all experiments, including those that require generalizing out-of-domain, the agents trained with augmentations perform considerably worse than those trained without augmentations (compare to Tables 2 and 3 in the main text). We hypothesize that since, unlike Dessì et al. (2021), we are using pre-trained visual models that already abstract away from low-level features of the images, there is not further benefit in applying augmentation during the communication-training phase. At the same time, this phase only tunes a single layer, which implies that the models do not have a lot of power to adapt to the augmentations. This ends up making the task harder without bringing about better abstractions skills than those already acquired during pre-training.
@@ Appendix I Visual-module to visual-module analysis
Appendix J 16-dimensional protocol generalization to unseen datasets
We repeat the experiments reported in from Table 18 with the 16-dimensional communication channel. Results show communication always remains much above random chance ($1/64\approx{}1.6\%$) in all three datasets. Nonetheless, the larger communication channel leads to better performance out of distribution, with up to 75% decrease of performance for the 16-dimensional messages in the setup were homogeneous pairs generalize from ImageNet1k to Places205.
@@
Appendix K Additional agent generalization with completely unseen models
@@
Appendix L Using CLIP as a visual module
@@
Appendix M Further SAE feature examples
@@
Appendix N Insights from communication errors and examples of ImageNet1k/Place205 message similarity
@@
Appendix O Discretization and other channel perturbations

### v7

@@ 4.4.2 Disentangling meaning with a Sparse Autoencoder
To investigate which high-level features are used in communication, we use dictionary learning, which has proven useful to identify interpretable features in multimodal models (Parekh et al., 2024), and more specifically Sparse AutoEncoders (SAEs) (Bricken et al., 2023; Cunningham et al., 2023; Rajamanoharan et al., 2024; Templeton et al., 2024). In our case, the SAE consists of single-layer encoders and decoders, with an added ReLu nonlinearity for the encoder. The encoder projects 64-dimensional messages to an expanded hidden space (we use a hidden space of $16\times{}64=1024$ dimensions). Training is done with a mean-square error reconstruction loss, and an L1 regularization loss to add sparsity (weight of the L1 term is set to 0.01). The sparsity component of the loss encourages storing meaningful information in individual dimensions. We train the SAE using messages corresponding to images from the communication training set (ImageNet1k, see section 3.3). Informally varying the SAE hyperparameters did not impact our general conclusions. We present SAE results for messages from a VGG 11 sender + ViT-B/16 receiver pair, which are representative of other cases.
@@
First and most importantly, we observe that the dimensions uncovered by the SAE correspond to understandable semantic features. In particular, we qualitatively assess this by looking, for each SAE dimension, at the images corresponding to the messages closest to that dimension. [^10] We were able to do this for all 1024 dimensions because many of them share the same neighbors. We observe that images clustered around a dimension share a common human-interpretable meaning. In Fig. 3, which contains the images closest to a representative SAE feature, we see different metallic grids. It is worth noting that they do not all share the same ImageNet1k class. The messages clustering around the same SAE vectors are abstract enough to overcome perceptual variation, with for example one of the images in Fig. 3 being photographed in night-time conditions (penultimate, first row), while the others all have good lighting conditions. We present additional representative examples in Appendix O (Fig. 11). Additional evidence for the semantic nature of the messages is provided in a qualitative error analysis in Appendix P.
@@
SAE features also inform us about referent generalization. We use the same sender and SAE trained on ImageNet1k to extract the k-nearest messages from the OOD datasets to each one-hot SAE vector. In general, the closest OOD messages to each SAE vector also form human-interpretable clusters, noisily sharing properties with the corresponding ImageNet1k training images. In the case of Places205, bottom right in Fig. 3, the images considered all contain an object with a metallic grid. For the more strongly OOD CelebA dataset, bottom left, pictures of celebrities tend to show either parallel lines, or in one case a metallic object. In general, OOD messages are futher away from one-hot SAE features than in-domain messages are, reflecting the fact that the corresponding images less closely match the properties captured by the features. [^11] On average, there are 56 ImageNet1k messages closer to a feature than the closest CelebA message. For Places205, which is less OOD, there are on average 7 ImageNet1K messages closer to a feature.
@@
From another perspective, instead of looking at which OOD messages are nearer pre-selected SAE vectors, we can ask which SAE vectors are closer to any message from an OOD dataset. Messages describing OOD images cluster around SAE vectors with properties relevant to the dataset. The CelebA messages tend to be closer to the SAE dimensions representing people, while Places205 messages are closer to SAE vectors corresponding to urban settings, or landscapes. [^12] To avoid pareidolia, we verified this claim in a small experiment. We prepared 30 sets of 10 ImageNet1k images, corresponding either to a feature close to Places205 images, or to one feature corresponding to CelebA images, or to neither. One of the authors then labeled them, without knowing their source, for whether they contained foregrounded faces (close to CelebA), depicted landscapes or urban settings (close to Places205), or neither. Performance was well above random chance. We provide an example of each in Fig. 12 of Appendix O.
@@ 4.4.3 Perturbation analysis
In a next set of analyses, we gain insight on the information encoded in the message by looking at how image perturbations affect communication (an analysis of direct perturbation of the communication channel itself is reported in Appendix Q). The way in which performance degrades in response to different perturbations should be informative about the image features that the communication protocol is relying on, and whether they are aligned with features that are intuitively crucial for our own high-level characterization of images. These experiments can also be seen as a further way to perform out-of-domain testing, where a protocol that was trained on non-augmented data is evaluated on perturbed images at test time.
The effects of the perturbations on an example image are shown on the right in Fig. 4. We perturb images from the ImageNet1k dataset that were seen by the agents during referential game training, where the receivers have perfect accuracy. In each experiment, we apply the same transformation to all images, both those seen by the sender and those seen by the receiver. Note that we test senders and receivers that have been trained with non-corrupted images, and only face the perturbations at test time.
Results for both the one-to-one and the population setups are reported in Fig. 4 (left). All ablations affect accuracy. However, the color ablations (transforming the image to grayscale and jittering colors) and resizing the image are having a much smaller impact on performance than adding Gaussian blur. This is in line with our intuition about how much these changes should affect the ability to decode the contents of an image. As the examples on the right in Fig. 4 show, color and size ablations don’t normally affect the human ability to discriminate the object depicted in the image, whereas blurring the image can easily make its contents difficult to decode for us as well.
We finally observe that all perturbation types affect one-to-one and population-trained agents in a similar manner, with population-trained agents being marginally more affected. [^13] If we continue training the population as we do in Appendix J, the gap between population-trained and one-to-one agents eventually disappears.
Figure 4: Accuracy after various image perturbations for the one-to-one and population setups (left). Gaussian blur was uniformly sampled within the [0.1,10] range. Mean accuracies and corresponding standard deviations are given across all 36 possible agent pairs. We add a small horizontal offset to the accuracy values for different perturbations to make them all visible. The horizontal lines connecting one-to-one and population accuracies for the same perturbation are meant to ease comparison between the two settings. Random baseline is 1.6% (randomly selecting one of the 64 images). Perturbation examples are provided on the right.
4.4.4 Specificity of population protocol
@@ 5 Conclusion
The networks can use the emergent protocol to communicate about referents that are only distinguishable at a more granular level than used during communication training; and they can communicate about referents coming from categories or even datasets that were not seen during pre-training nor in the communication protocol induction phase.
@@
Acknowledgments
We thank Agnieszka Słowik for early contributions to this work. We are also grateful to Thomas Brochhagen, Dan Dediu, Rahma Chaabouni, Emily Cheng, Valentino Maiorca, Luca Moschella, Clément Moulin Frier, the members of the COLT group and the TMLR editor and reviewers for useful feedback. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101019291). We also received funding from the Catalan government (AGAUR grant SGR 2021 00470). This paper reflects the authors’ view only, and the funding agencies are not responsible for any use that may be made of the information it contains.
@@ Appendix A The one-to-one referential game in pseudocode
Appendix B Using pre-trained vision modules without a communication layer
Our setup assumes that we need to connect networks that have already been pre-trained on visual tasks. An interesting question is to what extent the communication layer is adding something non-trivial to what is already encoded in the pre-trained visual representations of the networks. To assess this, we design a version of the communication game where the agents are pre-trained visual modules without any communication-specific fine-tuning.
Specifically, given two (distinct) pre-trained models that can predict image labels, we play the referential game as follows:
  • sender predicts label of target image;
  • receiver predicts label of all candidate images;
  • if sender and receiver labels of target differ, receiver picks a random image;
  • if labels coincide, the receiver picks a random image among those for which it has predicted the target label;
  • the game round is successful if the receiver picks the target image.
We experimented with all the models that were trained with the same label set (ViT-B/16, Inception, ResNet152, VGG11, Swin), using 10k images for each dataset. The results are in Table 6. The models can still communicate above chance level, but clearly communication layer training brings about a big improvement (compare to tables 2, 3, 4 and 5 above), confirming that this process is important to ensure successful referential communication across models. [^14] For technical reasons, we were not able to reproduce the Single-Class experiment for this setup. We expect however that performance would be extremely low for models that have been trained to predict the same label for all images presented in the same candidate set, as is the case with the Single-Class experiment.
Dataset Percentage accuracy
ImageNet (Val) 58.3 ± 21.3
ImageNet (OOD) 16.1 ± 4.3
Cifar100 9.3 ± 10.0
Places205 24.4 ± 13.0
CelebA 10.4 ± 5.2
Table 6: Percentage accuracy playing the communication game without a trained communication layer across various datasets, reported as mean ± standard deviation over all heterogeneous pairs.
@@ Appendix C Discrete experiments
In order to make sure that our negative results concerning discrete communication are not due to an unfortunate choice of hyperparameters, we repeated most experiments across a range of settings. We report here results for the more interesting population setup, but similar patterns are observed in one-to-one training. Table 9 shows that there is a small increase in accuracy when enlarging vocabulary size (compared to the results obtained with 256-dimensional messages reported in tables 7 and 8 above), but performance is still well below that of the continuous case even when using a vector as large as 8192 dimensions (compared to 16 or 64 dimensions in the continuous case). Note that using vocabulary sizes below 256 (not shown in the table) led to failure to converge to non-random accuracy.
@@
If using a discrete channel is considered a priority (for example, because it might lead to a more interpretable protocol), other discrete optimization techniques could also be explored, such as more advanced policy gradient methods (Francois-Lavet et al., 2018), or quantization approaches such as the one recently proposed by Carmeli et al. (2022) (although we tried something similar, with less-than-encouraging results, in Appendix Q).
@@ Appendix G Augmentations and hard distractors
Results for agents trained with data augmentations are presented in Table 16. In all experiments, including those that require generalizing out-of-domain, the agents trained with augmentations perform considerably worse than those trained without augmentations (compare to Tables 2 and 3 and 4 in the main text). We hypothesize that since, unlike Dessì et al. (2021), we are using pre-trained visual models that already abstract away from low-level features of the images, there is not further benefit in applying augmentation during the communication-training phase. At the same time, this phase only tunes a single layer, which implies that the models do not have a lot of power to adapt to the augmentations. This ends up making the task harder without bringing about better abstractions skills than those already acquired during pre-training.
@@ Appendix I Visual-module to visual-module analysis
Appendix J Longer population training
We let 64-dimensional population training run for 7 times the number of epochs (175 epochs) of one-to-one pair training, so that each single agent in the population is updated as often as the one-to-one homogeneous/heterogeneous agents. Results are in Table 19.
Dataset Accuracy
ImageNet (Val) 99.2 ± 1.1
ImageNet (OOD) 78.0 ± 15.3
ImageNet (Single Class) 43.6 ± 5.0
Places205 92.8 ± 6.7
CelebA 45.2 ± 24.8
Cifar100 37.6 ± 27.5
Table 19: Percentage accuracy for a population with longer training across all different datasets. Values are reported as mean ± standard deviation across models in the population.
While performance appeared to plateau, if we keep training, we observe instead non-negligible improvements across experiments (compare to tables 2, 3, 4 and 5 above). Still, in all cases except for the in-domain ImageNet set, performance still lags behind that of the heterogeneous pairs, that were trained in 1/7 of the time required to bring population performance at the level reported here.
Appendix K 16-dimensional protocol generalization to unseen datasets
We repeat the experiments reported in Table 5 with the 16-dimensional communication channel. The results (in Table 20) show that communication always remains much above random chance ($1/64\approx{}1.6\%$) in all three datasets. Nonetheless, the larger communication channel leads to better performance out of distribution, with up to 75% decrease of performance for the 16-dimensional messages in the setup where homogeneous pairs generalize from ImageNet1k to Places205.
@@
Appendix L Sequential population training
In the main paper, we train population agents in parallel, by pairing a random sender and receiver from the whole population at each step. We explore here the alternative in which we initially train a pair of agents for a number of epochs, and then progressively let the community grow one agent at a time.
Two important hyperparameters of sequential population training are agent order and training schedule, namely how many epochs are given to each new agent before introducing the next. Concerning agent order, we tried 10 random sequences, and we report results averaged across them. Concerning training schedule, we considered two setups. In the basic experiment, we trained the first pair for 2 epochs, and then trained for 2 epochs more after adding each new agent (sender or receiver), for a total of 30 epochs. This is comparable to the total amount of training of the population used for the main experiments (25 epochs). In the long experiment, we train the initial pair for 7 epochs, and then, for each new agent we introduce, we train for 7 more epochs. This is based on the observation that pairwise convergence tends to take 7 epochs or less. In this case, we are running a total of 105 epochs, closer to the training length of the experiment we report in Appendix J above (175 epochs). For both setups, we adopt a 64-dimensional communication vector.
Dataset Basic Long
ImageNet (Val) 94.5 ± 9.2 98.1 ± 3.1
ImageNet (OOD) 45.0 ± 22.7 72.1 ± 18.2
ImageNet (Single Class) 32.6 ± 11.1 38.4 ± 8.8
Places205 63.6 ± 26.0 91.1 ± 8.3
CelebA 33.4 ± 23.1 40.0 ± 25.2
Cifar100 27.4 ± 25.3 35.6 ± 27.7
Table 21: Percentage accuracy for two sequential training lengths on all datasets. Values are reported as mean ± standard deviation across sequence orders.
In both setups, sequential training leads to results that are inferior to those of the corresponding parallel-training experiments: cf. tables 2, 3, 4 and 5 in the main text for the basic results and Table 19 in Appendix J for the long results.
We conjecture that the problem with sequential training is that the first agents to be trained will disproportionately affect the emergent protocol, making it harder for later agents to adapt. Given also that sequential training requires more hyperparameters to tune, the takeaway here is that parallel training is the better option.
Appendix M Additional agent generalization with completely unseen models
@@
Appendix N Using CLIP as a visual module
@@
Appendix O Further SAE feature examples
@@
Appendix P Insights from communication errors and examples of ImageNet1k/Place205 message similarity
@@
Appendix Q Discretization and other channel perturbations